Write the slope intercept form of the equation of the line through the given point, with the given slope.
(0), (-5), slope = 0

Answer by hailee.devoll(2)

y=mx+b plug the points in -5=0(0)+b so b=-5 so plugg b back in and you get for your answer is y=-5